OpenStudy (agl202):

Medal Fan for best answer!!! Line AB contains points A (8, -4) and B (1, -5). The slope of line AB is -7 -1/7 1/7 7

OpenStudy (anonymous):

Plug them into the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-5+4}{1-8}\) Simplify, and see if the number is positive or negative.

OpenStudy (anonymous):

1/7

OpenStudy (anonymous):

@hamidic Don't give answers, it's against the code of conduct.

OpenStudy (anonymous):

Add -5 + 4 and subtract 1 - 8, what do you get? @Agl202

OpenStudy (anonymous):

@iGreen. wat's the code of conduct ?! ... sorry it's first tie i'm here science 2 years